A heat engine takes in 1500 J of heat at 2000 K and rejects 950 J of heat at 1000 K. Find the (a)   maximum efficiency of this engine and (b) its actual efficiency.


1
Expert's answer
2021-12-14T09:27:52-0500

(a) "\eta=(T_1-T_2)/T_1=(2000-1000)/2000=0.5"


(b) "\eta=(Q_1-Q_2)/Q_1=(1500-950)/1500=0.37"